OpenStudy (anonymous):

Need a little help Solve the system using elimination 3x + 5y = −18 4x − 10y = −24

OpenStudy (shylovesyou11):

3x+5y=-18 call this equation 1 4x-10y=-24 call this equation 2 This is quite easy, let's just multiply equation 1 through by 2. 6x + 10y = -36 now let's add this to equation 2 4x-10y=-24 10x = -60 divide by 10 x = -6 now substitute back to get y. 4x-10y=-24 4(-6) -10y = -24 -24 -10y = -24 -10y = -24+24 -10y = 0 y=0 =)

OpenStudy (anonymous):

Oh thanks I was confused on whether or not to multiply by 2 or not!!

OpenStudy (shylovesyou11):

You're welcome! ^_^